# Easy conditioning far beyond Gaussian

### v3

@@ 2.2.1 Extension to finite mixtures
Let us first define a “mixing” operator which takes as input a parameterized family of multivariate pdfs with a fixed dimension $d\geq 1$ and returns the set of probabilities densities that are finite mixtures of elements of this family.
Consider again a parametrized family $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\Theta\in\Theta^{(d)}\}$ of pdfs. We define the mixing operator as:
@@
We now present the crucial results that extend the concepts of stability by marginalization and conditioning to finite mixtures.
@@
Proof: Let consider $f$ in $\text{Mix}(\mathcal{F})$. There exists $d\in\{1,\ldots,D\}$, $K\geq 1$, $\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)}$, $\alpha_{1},\ldots,\alpha_{K}\geq 0~{}\text{s.t}~{}\sum_{k=1}^{K}\alpha_{k}=1$ such that:
@@
Since $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, then for every $k\in\{1,\ldots,K\},~{}f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\in\mathcal{F}$ and the conditional distribution $f_{\mathbf{i}|\mathbf{j}}$ is a mixture of elements of $\mathcal{F}$ with weights $\tilde{\alpha}_{k}$ so we can conclude that $f_{\mathbf{i}|\mathbf{j}}\in\text{Mix}(\mathcal{F})$. The proof for marginalization is analogous. $\square$
This result can similarly be applied to cdfs. Note that if we can compute analytically the conditioned mixing weights $\tilde{\alpha}_{k}$ and if we can sample from the conditional distributions of each component $f_{\mathbf{i}|\mathbf{j}}(\cdot;\theta_{k})$, we can produce conditional samples of a mixture of distributions.
@@ 2.2.2 Extension to transformations
We define, similarly as in the previous section for pdfs, parametric families of cdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$ as $\mathcal{F}^{d}=\left\{F^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\right\}=\left\{F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$ and trans-dimensional families as $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D\leq+\infty$.
For fixed $d\geq 1$ and $\mathcal{X}^{\prime}_{i},~{}i=1,\ldots,d$ measure spaces, we define the class of functions:
@@
where $\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X^{\prime}}_{i})$ is the class of increasing $\mathcal{C}^{1}$ diffeomorphisms from $\mathcal{X}_{i}$ to $\mathcal{X}^{\prime}_{i}$. We introduce a “push-forward” operator, which takes as input an element of a parametric family of cdfs and returns every possible cdfs obtained by applying marginal $\mathcal{C}^{1}_{\uparrow}$ diffeomorphisms to the components, by:
@@
Note that elements of $\Xi^{(d)}\#F^{(d)}_{\theta}$ are probability distributions on $\prod_{i=1}^{d}\mathcal{X^{\prime}}_{i}$ and that we restricted to differentiable bijections such that the obtained probability distributions are still absolutely continuous.
@@
If we define the class of functions with increasing marginal $\mathcal{C}^{1}$ diffeomorphisms as
$\Xi=\bigcup\limits_{d=1}^{D}\Xi^{d}$
we can now apply the "push-forward" operator to a trans-dimensional family of cdfs $\mathcal{F}$ by:
@@
We call the family of probability distributions $\Xi\#\mathcal{F}$ the push-forward family of $\mathcal{F}$-distributions. In the next section, we will relate this concept to copulas and meta families of distributions. At this point, we present the results that extend the stability of marginalization and conditioning to transformations.
@@
Proof: Let $F\in\Xi\#\mathcal{F}$. There exists $d\in\{1,\ldots,D\},~{}\theta\in\Theta^{(d)},~{}\xi\in\Xi^{d}$ such that:
$\forall(x_{1},\ldots,x_{d})\in\prod_{k=1}^{d}\mathcal{X}_{k},~{}F(x_{1},\ldots,x_{d})=(F^{(d)}\circ\xi^{-1})(x_{1},\ldots,x_{d};\theta)=F^{(d)}(\xi_{1}^{-1}(x_{1}),\ldots,\xi_{d}^{-1}(x_{d});\theta)$
@@
with $z_{i_{k}}=\xi^{-1}_{i_{k}}(x_{i_{k}}),k=1,\ldots,\ell$ and $z_{j_{k}}=\xi^{-1}_{j_{k}}(x_{j_{k}}),k=1,\ldots,m$. Since $\mathcal{F}$ is stable by conditioning, then $F^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot;\theta)$ belongs to to $\mathcal{F}$. The proof is similar for marginalization. $\square$
@@
Note that we have only considered univariate transformations. However, for a fixed set of conditioning indices, it is worth mentioning that by following the same reasoning as in the proof, we can extend the property to blockwise bijective transformations. This goes beyond the scope of the present paper.
3 Methodology for Estimating Conditional Distributions
As highlighted in the introduction, our focus is on estimating the conditional distributions of a (multivariate) response variable. This section delineates our methodology.
Assume we want to condition on an given set of variables. Without loss of generality, we can rearrange the components of a random variable $\mathbf{X}$ in two groups of variables $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ such that the first group $\mathbf{X_{1}}$ contains the variables that we want to estimate conditional on the second group of variables $\mathbf{X}_{2}$. By assuming that we can fit our models on a previously acquired sample $(\mathbf{X_{1}},\mathbf{X_{2}})$ taking values on a continuous scale, and given a new an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}$, we derive a procedure to estimate the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x}_{2}$.
We consider the scenario of real-valued random variables, corresponding to $\prod_{k=1}^{d}\mathcal{X}_{k}=\mathbb{R}^{d}$ and we focus exclusively on absolutely continuous random variables.
Our generative procedure to estimate a conditional distribution can be divided in two main steps:
  • estimate the joint distribution by a parametric form.
  • perform conditioning of the estimated joint by using analytical formulas.
In modeling the joint distribution, much of the existing literature focuses on the Gaussian case due to the simplicity of its analytical formula for conditional distributions. However, multivariate Gaussian distributions have notable limitations, as they often fail to adequately model skewed, heavy-tailed, or multi-modal data.
Alternative multivariate distributions, such as Student’s t-distributions [11] or skew-normal distributions [3], offer more expressiveness and are stable by conditioning, see appendix for details. Yet, these distributions also fall short in scenarios requiring the modeling of multi-modal data and are confined to have marginal distributions of the same family along each axis.
By leveraging the extension of stability by conditioning to mixtures and transformations discussed in the preceding section, we achieve considerably greater flexibility in modeling the joint distribution. The extension to transformations suggests that we can identify a latent space where stability by conditioning holds. To develop a fast and generic approach for estimating these conditional distributions, our interest lies particularly in families of distributions for which conditional distributions are known in closed form.
By using univariate transformations, we can model distributions with arbitrary marginals. Specifically, if we consider the univariate transformations to be the marginal cumulative distribution functions, we obtain a space with uniform marginals and can subsequently apply copula theory, as detailed in the next section.
3.1 Estimate the joint multivariate distribution with copulas
@@
From a $d$-dimensional sample $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ of independently drawn realizations, we want to estimate the distribution of the underlying random vector $\mathbf{X}$ taking values in $\mathbb{R}^{d}$. We will use a two-steps procedure based on copulas.
@@
Copulas provides a versatile framework for multivariate analysis. Following Sklar’s theorem [22], for any distribution function $F$ with continuous margins $F_{j},~{}j=1,\ldots,d$, there exists a unique copula $C$ such that
@@
There exist several parametric families of copulas. In particular, let’s introduce the notion of “implicit copula” which refers to the copula implicit in the multivariate distribution of a continuous random vector.
@@
where $z_{j}=F_{j}^{-1}(u_{j}),~{}j\in\{1,\ldots,d\}$, we say that $\mathbf{z}=(z_{1},\ldots,z_{d})$ is the quantile transform of $\mathbf{u}$.
@@
However, in practice we don’t know the marginal distributions and we need to estimate them from the data. Parametric or non-parametric methods can be used to estimate the marginal distributions. Parametric estimation of the marginals is used in the inference for margins (IFM) method [15], while non-parametric estimation using the empirical distribution function is part of the canonical maximum likelihood estimation (CMLE) method [13], which has been proven to be asymptotically consistent.
@@
A $d$-variate meta distribution is constructed by linking a $d$-variate implicit copula with $d$ arbitrary marginal cdfs. For example, a meta Gaussian distribution has a Gaussian copula with arbitrary marginals. Meta elliptical distributions have been studied [12].
More formally, for a family of distributions $\mathcal{F}^{(d)}$ in dimension $d$ such as Gaussian or Student $t$-distributions we define its meta family as :
$\text{Meta}(\mathcal{F}^{(d)})=\left\{F(\cdot)=(C_{\theta}\circ\tilde{F})(\cdot),~{}\theta\in\Theta^{(d)},~{}\tilde{F}\in\mathcal{M}^{d}\right\}$
where $\left\{C_{\theta},~{}\theta\in\Theta^{(d)}\right\}$ is a family of parametric copulas in dimension $d$ such that $\forall\theta\in\Theta^{(d)},~{}F_{\theta}\in\mathcal{F}^{(d)}~{}\text{has copula}~{}C_{\theta}$ and :
$\mathcal{M}^{d}=\left\{F=\begin{pmatrix}F_{1}\\ \vdots\\ F_{d}\end{pmatrix},~{}F_{1},\ldots,F_{d}~{}\text{absolutely continuous and strictly increasing cdfs}\right\}$
And for a trans-dimensional family $\mathcal{F}$ defined as in the last section, we get:
$\text{Meta}(\mathcal{F})=\bigcup\limits_{d=1}^{D}\text{Meta}(\mathcal{F}^{(d)})$
We can remark that for a given family $\mathcal{F}$ its push-forward family includes its meta family. The proof showing that stability by conditioning of a given trans-dimensional family $\mathcal{F}$ implies the stability by conditioning of its meta family, closely resembles that of Theorem 2.2.
In our procedure, we assume that the joint distribution of the data is a meta distribution with a parametric copula from a well known family. In particular, we choose families of distributions for which are stable by conditioning. For example, the conditional distributions of a meta Gaussian distributions are also meta Gaussian.
We should mention that the stability of meta distributions corresponds to the stability of copulas. Our research focuses on parametric families of copulas for which the conditional copulas belongs to the same parametric family.
3.2 Conditioning in a latent space
Once we have estimated the joint distribution, the second step of our method is to estimate the conditional distribution (get conditional samples) given a new observation of the covariates. Since we assume that the data follow a "meta stable" distribution in order to perform exact conditioning, we will be able to find a latent space in which we have stability by conditioning. In practice, we will focus on distributions for which there is an analytical formula for the conditional distributions.
Copulas transform the marginals to the uniform space $\mathcal{U}$ because it is usually easier to capture multivariate dependence using $C$ on $[0,1]^{d}$. We go one step further by transforming the marginals to a latent space $\mathcal{Z}$ [23], in which the marginals are the margins of an implicit multivariate distribution driving the copula. The two consecutive transformations are explained in the following scheme:
where $\Psi_{i}(\cdot;\hat{\theta})$ represents the cumulative distribution function of the $i$-th marginal of an implicit multivariate distribution linked to the copula parameters $\hat{\theta}$ estimated in the last section. In the latent space $\mathcal{Z}$, the samples follow this implicit multivariate distribution.
@@ 3.2 Conditioning in a latent space
Algorithm 1 Estimate the cdf of a conditional distribution
0:  $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ a sample from $\mathbf{X}\in\mathbb{R}^{d}$, $\ell$ the dimension of the response, a new point $\mathbf{x}^{(new)}_{\ell+1:d}=(x^{(new)}_{\ell+1},\ldots,x^{(new)}_{d})$.
@@
    $\hat{F}_{\mathbf{X}_{1}|\mathbf{X}_{2}=\mathbf{x}^{(new)}}(x_{1},\ldots,x_{\ell}|\mathbf{z}_{\ell+1:d}^{(new)})=\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}=\mathbf{z}^{(new)}_{\ell+1:d}}(\hat{F}_{1}^{-1}\circ\Psi_{1}(z_{1}^{(new)}),\ldots,\hat{F}_{\ell}^{-1}\circ\Psi_{\ell}(z_{\ell}^{(new)})|\mathbf{z}_{\ell+1:d}^{(new)})$

### v4

@@ 2.2.1 Extension to finite mixtures
Consider again a parametrized family $\mathcal{F}^{(d)}=\{f^{(d)}(\cdot;\theta);\Theta\in\Theta^{(d)}\}$ of pdfs. We define
@@
We now present results that extend the concepts of stability by marginalization and conditioning to finite mixtures.
@@
Proof of Theorem 2.1.
Let consider $f$ in $\text{Mix}(\mathcal{F})$. There exists $d\in\{1,\ldots,D\}$, $K\geq 1$, $\theta_{1},\ldots,\theta_{K}\in\Theta^{(d)}$, $\alpha_{1},\ldots,\alpha_{K}\geq 0~\text{s.t}~\sum_{k=1}^{K}\alpha_{k}=1$ such that:
@@
Since $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, then for every $k\in\{1,\ldots,K\},~f^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot|(x_{j_{1}},\ldots,x_{j_{m}});\theta_{k})\in\mathcal{F}$ and the conditional distribution $f_{\mathbf{i}|\mathbf{j}}$ is a mixture of elements of $\mathcal{F}$ with weights $\tilde{\alpha}_{k}$ so we can conclude that $f_{\mathbf{i}|\mathbf{j}}\in\text{Mix}(\mathcal{F})$. The proof for marginalization is analogous. ∎
This result can similarly be applied to cdfs. Note that if we can compute analytically the conditioned mixing weights $\tilde{\alpha}_{k}$ and if we can sample from the conditional distributions of each component $f_{\mathbf{i}|\mathbf{j}}(\cdot;\theta_{k})$, we can produce samples from the conditionals of the overarching mixture of distributions. Although the authors are not aware of prior formulations of Theorem 2.1, it its worth remarking that the formula for the conditional distribution of a mixture of distributions has in fact been utilized in diverse contexts [gilardi2002conditional, spycher2009multivariate].
@@ 2.2.2 Extension to transformations
We now work with parametric families of cdfs on $\prod_{k=1}^{d}\mathcal{X}_{k}$, similarly denoted denoted $\mathcal{F}^{d}=\left\{F^{(d)}(\cdot;\theta);\theta\in\Theta^{(d)}\right\}=\left\{F^{(d)}_{\theta};\theta\in\Theta^{(d)}\right\}$ as done with pdfs in the previous section. Similarly, we denote resulting trans-dimensional families as $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D\leq+\infty$.
For fixed $d\geq 1$ and $\mathcal{X}^{\prime}_{i}(i=1,\ldots,d$) measure spaces, we further define
@@
where $\mathcal{C}^{1}_{\uparrow}(\mathcal{X}_{i},\mathcal{X^{\prime}}_{i})$ is the class of increasing $\mathcal{C}^{1}$ diffeomorphisms from $\mathcal{X}_{i}$ to $\mathcal{X}^{\prime}_{i}$. We introduce a “pushforward” operator, which takes as input an element of a parametric family of cdfs and returns all possible cdfs obtained by applying marginal $\mathcal{C}^{1}_{\uparrow}$ diffeomorphisms to the components, by:
@@
Note that elements of $\Xi^{(d)}\#F^{(d)}_{\theta}$ are probability distributions on $\prod_{i=1}^{d}\mathcal{X^{\prime}}_{i}$ and that the obtained probability distributions are still absolutely continuous.
@@
Defining finally the class of functions with increasing marginal $\mathcal{C}^{1}$ diffeomorphisms as
$\Xi=\bigcup\limits_{d=1}^{D}\Xi^{d},$
we can now apply the pushforward operator to a trans-dimensional family of cdfs $\mathcal{F}$ by:
@@
We call the trans-dimensional family $\Xi\#\mathcal{F}$ the pushforward of $\mathcal{F}$. In the next section, we relate this concept to copulas and meta families of distributions. At this point, we extend the stability of marginalization and conditioning results to transformations.
@@
Proof of Theorem 2.2.
: Let $F\in\Xi\#\mathcal{F}$. There exists $d\in\{1,\ldots,D\},~\theta\in\Theta^{(d)},~\xi\in\Xi^{d}$ such that $\forall(x_{1},\ldots,x_{d})\in\prod_{k=1}^{d}\mathcal{X}_{k}$:
$F(x_{1},\ldots,x_{d})=(F^{(d)}\circ\xi^{-1})(x_{1},\ldots,x_{d};\theta)=F^{(d)}(\xi_{1}^{-1}(x_{1}),\ldots,\xi_{d}^{-1}(x_{d});\theta)$
@@
with $z_{i_{k}}=\xi^{-1}_{i_{k}}(x_{i_{k}}),k=1,\ldots,\ell$ and $z_{j_{k}}=\xi^{-1}_{j_{k}}(x_{j_{k}}),k=1,\ldots,m$. Since $\mathcal{F}$ is stable by conditioning, then $F^{(d)}_{\mathbf{i}|\mathbf{j}}(\cdot;\theta)$ belongs to to $\mathcal{F}$. The proof is similar for marginalization. ∎
@@
Note that we have only considered univariate transformations. However, for a fixed set of conditioning indices, it is worth mentioning that by following the same reasoning as in the proof, we can extend the property to blockwise bijective transformations.
3 Easy conditioning with fitted stable by conditioning distributions
We now focus on the task of conditioning an absolutely continuous random vector $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ taking values in $\prod_{k=1}^{d}\mathcal{X}_{k}=\mathbb{R}^{d}$ on the second group of variables $\mathbf{X}_{2}$. On our approach is to first fit a joint (easy to condition) distribution to previously acquired sample $(\mathbf{x_{1i}},\mathbf{x_{2i}})_{i=1}^{n}$, and then, given any new observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}$, to leverage our results to estimate and/or sample from the conditional distribution of $\mathbf{X}_{1}$ given that $\mathbf{X}_{2}=\mathbf{x}_{2}$. Our generative procedure to estimate a conditional distribution thus consists in two main steps:
  • Fit an “easy to condition” joint distribution to a sample from $\mathbf{X}$
  • perform conditioning of the estimated joint by leveraging analytical formulas.
By using univariate transformations, we can model distributions with arbitrary (absolutely continuous) marginals. Specifically, if we consider the univariate transformations to be the marginal cumulative distribution functions, we obtain a space with uniform marginals and can subsequently apply copula theory, as detailed in the next section. In order to perform easy conditioning, we favor copula models corresponding to distributions stable by conditioning. Gaussian copulas, in particular, offer analytical formula for conditional distributions. However, they rely on latent multivariate Gaussian distributions, which are known to have notable limitations in terms of adequately model skewed, heavy-tailed, or multi-modal data. Alternative multivariate distributions, such as Student’s t-distributions [ding2016conditional] or skew-normal distributions [azzalini2013skew], offer more expressiveness and are stable by conditioning (see appendix for details). Resulting copula models can be used in our framework as they enable easy conditioning too. Yet, these distributions also fall short in scenarios requiring the modeling of multi-modal data and are confined to have marginal distributions of the same family along each axis. An important aspect of our the practical approach presented next is to allow capturing multi-modality by leveraging the extension of stability by conditioning to mixtures. The workhorse is to appeal to copulas of mixtures of multivariate distributions, where each individual component multivariate distribution possesses known conditional distributions.
3.1 Fitting joint distributions in terms of marginals and copulas
@@
From a $d$-dimensional sample $\mathcal{D}=(x_{1i},\ldots,x_{di})_{i=1}^{n}$ of independently drawn realizations, we want to estimate the distribution of the underlying random vector $\mathbf{X}$ taking values in $\mathbb{R}^{d}$. We use here a two-step procedure based on copulas.
@@
Copulas provides a versatile framework for multivariate analysis. Following Sklar’s theorem [sklar1959fonctions], for any distribution function $F$ with continuous margins $F_{j},~j=1,\ldots,d$, there exists a unique copula $C$ such that
@@
There exist several parametric families of copulas. In particular, for a given absolutely continuous multivariate distribution, we denote by “implicit copula” the underlying the latter.
@@
where $z_{j}=F_{j}^{-1}(u_{j}),~j\in\{1,\ldots,d\}$, we say that $\mathbf{z}=(z_{1},\ldots,z_{d})$ is the quantile transform of $\mathbf{u}$, and $f$ and $f_{i}$ denote the densities associated with $F$ and $F_{i}$, respectively, for all $i\in{1,\ldots,d}$.
@@
However, in practice the marginal distributions are typically unknown, and one needs to estimate them from the data. Parametric or non-parametric methods can be used to estimate the marginal distributions. Parametric estimation of the marginals is used in the inference for margins (IFM) method, while non-parametric estimation using the empirical distribution function is part of the pseudo likelihood estimator used in [genest1995semiparametric], which has been proven to be allow for asymptotically consistency in the estimation of dependence parameters.
@@
In our procedure, we assume that the joint distribution of the data is a meta distribution. A $d$-variate meta distribution is constructed by linking a $d$-variate implicit copula with $d$ arbitrary marginal cdfs. For example, a meta Gaussian distribution utilizes a Gaussian copula paired with arbitrary marginals.
In this study, we work with parametric copulas implicit in a family of distributions stable by conditioning. Consequently, as derived from Theorem 2.2, when considering strictly increasing marginal cdfs as a case of $\mathcal{C}_{\uparrow}^{1}(\mathbb{R},\mathbb{R)}$, the meta distributions we consider also exhibit stability by conditioning. For practical applications, our focus will be on copulas implicit in families of distributions for which analytical computation of conditional distributions is possible.
3.2 Conditioning via latent space
Once we have estimated the joint distribution, we estimate the conditional distribution (get conditional samples) given a new observation of the covariates. Since we assume that the data follow a "meta stable" distribution in order to perform easy conditioning, we are able to find a latent space in which we have stability by conditioning. In practice, we will focus on distributions for which there is an analytical formula for the conditional distributions.
Copulas transport the problem from the original space $\mathcal{X}$ to uniform space $\mathcal{U}=[0,1]^{d}$. We go one step further by transforming the marginals to a latent space $\mathcal{Z}$ [smith2023implicit], in which the marginals are the those of an implicit multivariate distribution driving the copula. The two consecutive transformations are illustrated in the following scheme:
[…]
@@ 3.2 Conditioning via latent space
Algorithm 1 Estimation of conditional CDF via latent space
[…]
@@
    $\hat{F}_{\mathbf{X}_{1}|\mathbf{X}_{2}}\left(x_{1},\ldots,x_{\ell}\big|\mathbf{x}_{\ell+1:d}^{(new)}\right)=\hat{F}_{\mathbf{Z}_{1}|\mathbf{Z}_{2}}\left(\hat{F}_{1}^{-1}\circ\Psi_{1}(z_{1}^{(new)}),\ldots,\hat{F}_{\ell}^{-1}\circ\Psi_{\ell}(z_{\ell}^{(new)})\big|\mathbf{z}_{\ell+1:d}^{(new)}\right)$